OpenStudy (anonymous):

f(x)=6x-6 and g(x)=1/6(x+6) What is f(g(x))? Any given values of x≠? What is g(x(x))? Any given values of x≠? And are the functions of f and g inverses of each other?

OpenStudy (anonymous):

i assume this is \[g(x)=\frac{x+6}{6}\]

OpenStudy (anonymous):

so f and g are inverses because \[f(g(x))=6\times (\frac{x+6}{6})-6=x+6-6=x\]

OpenStudy (anonymous):

there are no restrictions on x because both f and g are polynomials

OpenStudy (anonymous):

\[g(g(x))=g(\frac{x+6}{6})=\frac{\frac{x+6}{6}+6}{6}\]

OpenStudy (anonymous):

a little algebra (muliply top and bottom by 6) gives \[\frac{x+6+36}{36}=\frac{x+42}{6}\]